A Space & astronomy forum. SpaceBanter.com

Go Back   Home » SpaceBanter.com forum » Space Science » Space Station
Site Map Home Authors List Search Today's Posts Mark Forums Read Web Partners

ISS On-Orbit Status, 15-12-2003



 
 
Thread Tools Display Modes
  #1  
Old December 16th 03, 10:12 AM
Jacques van Oene
external usenet poster
 
Posts: n/a
Default ISS On-Orbit Status, 15-12-2003

ISS On-Orbit Status 15 Dec 2003

All ISS systems continue to function nominally, except as noted previously
or below. This is Week 8 for Expedition 8.

TsUP/Moscow reported "no issues" with this morning's first part of the R&R
(removal and replacement) of the heat exchanger (BTA) of the Russian
segment's SKV-2 air conditioner. FE Alexander Kaleri performed drainage and
venting of the Khladon (Freon-218) to space on time and installed the BTA on
the SKV-2's condenser unit (BK). Completing hydraulic and electric
connections remains to be done tomorrow, followed by closeout ops and
checkout activities. [Changing the SKV-2's heat exchanger became necessary
because of inefficient performance of the old BTA, which equated to lack of
redundancy for the operating SKV-1. The R&R activity, similar to an activity
performed by Exp. 4's Yuri Onufrienko & Carl Walz on SKV-1 on 4/3/2002, was
originally planned for Increment 7 but was cancelled since the safety
certificate for the necessary venting of liquid Khladon, a toxic Level 1
hazard, wasn't available yet. Today's ops required activation of two pumps
(KOB1 & KOB2) of the Russian SOTR thermal control system after inhibiting
transfer to Survival mode per SOTR operating criteria, deactivation of
Elektron and SKV-1 during the Khladon venting, and returning all systems to
the initial configuration afterwards. The venting was started on TsUP Go
during an RGS (Russian ground sites) pass and conducted during the sunlit
portion of the orbit for observation.]

Elektron Update: As reported, the Elektron water electrolysis unit in the
Service Module (SM) has experienced several shutdowns recently, three over
the past weekend alone. The O2 generator has now been switched to its lowest
power level (16 amps), which may allow longer operating periods between pump
shutdowns. At this load setting, ppO2 (oxygen partial pressure) drops by
0.5-0.6 mmHg per day. TsUP estimates that this mode, supplemented by the O2
supply stored in Progress 12P, will provide two months of oxygen, enough to
last until Progress 13P brings new O2. [The pump shutdowns (primary pump
first, followed after ~20 min. by the backup unit) are thought to be caused
by pump impeller cavitation of air bubbles in the water, pointing to
probable malfunction of the membrane in the air/liquid separation component
of the Liquid Unit (BZh), which, as final solution, would have to be
replaced. The unit measures 30 cm in diameter, 1.1 m in height and 90 kg of
mass. The current redesigned BZh was installed last year.]

12P O2 Repress Update: Next repress of the stack with Progress 12P oxygen is
scheduled for 12/17 (Wednesday). The U.S. MCA (major constituent analyzer)
will be activated ~12 hrs before and calibrated ~6 hrs before the repress.
The 12P O2 Repress Plan has been worked out with Moscow for best utilization
of the 21 kg (12 days) of oxygen supply stored in 12P. ppO2 is measured by
the SM GA (gas analyzer), supported by the U.S. CSA-CP (compound specific
analyzer-combustion products). For repress planning, Moscow uses the GA
data, which are considered more accurate for this purpose. 12P represses are
performed when GA or CSA-CP ppO2 data drop below 146 mmHg (allowable
minimum: 120 mmHg). [Total onboard O2 supply for the crew is currently ~310
kg. Redline, with 45-day skip cycle, is at 115 kg (135 SFOGs, solid fuel
oxygen generators; usage: 2 units/day/2 crew).]

After transferring the newly assembled third PCS (portable computer system)
laptop from the Airlock (A/L), CDR Michael Foale closed the Node starboard
hatch and began the scheduled leak check of two hatch valves. "Quest" will
become accessible to the crew again tomorrow after A/L isolation and partial
depressurizing have been reversed. [The leak check of the MPEV (manual
pressure equalization valve) and IMV (intermodular ventilation) valve had
not been performed after the valves were removed, then reinstalled, when the
Exp 6/7 crews were unable to reopen the hatch after closing it as a handover
activity. Today's activity is to verify good seals on these valves. The CDR
used the depress pump to lower the A/L pressure to ~560 mm Hg (10.8 psia)
from 760 mmHg regular, while the ground increased the pressure of the rest
of the stack to 770 mm Hg (14.9 psia). Mike also installed a cap over the
depress pump feed-through line in the Node, which was found on 7A to have
leakage from A/L to Node that would invalidate the leak check. MCC-H
activated the A/L CCAA (common cabin air assembly) air conditioner during
the leak check.]

The planned periodic O-OHA (on-orbit hearing assessment) audiogram test for
the CDR was deferred due to lack of time on today's schedule.

Mike Foale activated the MSG (Microgravity Science Glovebox), then removed
the PFMI (Pore Formation &Mobility Investigation) sample from its thermal
chamber, leaving it in the MSG work volume for future troubleshooting of the
recent circuit breaker trip due to an overcurrent spike (6 amps) from the
PFMI. [Foale and the Principal Investigator (PI) held a teleconference,
supported by good photographs obtained by Mike after the sample removal.
Thirteen experiments have been successfully conducted and eight more are
planned for Increment 8. The micro-G PFMI investigation, with its capability
to control and visualize melting and solidification over a wide range of
temperature gradient and translation parameters, is promoting our knowledge
of bubble/porosity movement, solid/liquid interface dynamics, and
solidification phenomena.]

Prior to the SKV-2 R&R, Foale deployed passive FMK (formaldehyde monitoring
kit) collection badges in Lab and SM, to catch any atmospheric formaldehyde
on a sampling substrate.

Sasha Kaleri completed the periodic servicing task of changing out ASU
toilet system equipment, then checked out the ASU. [This involved a
receptacle (MP) and a filter insert (F-V). The old units were stowed for
disposal.]

A preliminary Progress 12P pre-pack list of 45 items was uplinked to help
the crew in stowing excess equipment and waste in the resupply ship, which
has already received liquid waste (urine) in its two Rodnik water tanks.
[Medical support system (CMO) trash is to be transferred from mesh bags to
soft trash bags (KBO-M), of which nine are to be placed in 12P.]

The crew worked out according to their regular daily physical exercise
program of 2.5 hrs on the TVIS treadmill, RED exerciser, CEVIS cycle
ergometer and VELO bike (with load trainer). Mike performed the weekly
maintenance of the treadmill and conducted a special inspection of the newly
installed wire ropes.

The FE prepared the daily IMS (inventory management system) "delta" file for
automatic export/import to update the database; he also performed the daily
routine maintenance of the SOZh life support system.

Mike Foale had his weekly crew conference with the JSC Astronaut Office.

At 9:50am EST, the CDR reconfigured the TV hardware setup for two
back-to-back interactive TV interview events starting at 10:10am, with the
British GQ Magazine and BBC "Look East" Radio Network.

Weekly Science Update (Expedition Eight -- 6th):

GASMAP: "Thank you" to Mike for performing the GASMAP health check on 12/12.
Looking forward to hearing from the crew again next month.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound: Planned.

Hand Posture Analyzer (HPA): Next session is planned for mid-increment, and
a third late in the increment.

In-Space Soldering Investigation (ISSE): Planned.

Pore Formation and Mobility Investigation (PFMI): Ground specialists are
currently examining the downlinked data to better understand the cause of
the circuit breaker trip on 12/8. PFMI ground hardware was returned from JSC
on 12/11 from crew training so now some testing can be done with it. One
possible cause is the drive motor firmware. The gradient stage and camera
motors are "smart" motors with their own firmware and have proven to be
susceptible to Single Event Upsets before. Most recently, 8/20/03, the drive
motor would not properly accept commands to translate. The good news is that
if this is the cause of the CB-1 trip, a simple recycling of power resets
the firmware. Another area being looked at are the motor gears. There may be
some kind of debris or obstruction causing jamming. Today's discussion with
Mike was very helpful. Space Acceleration Measurement System (SAMS): SAMS
operations are back to normal after the system was restored tonominal
configuration. Microgravity Acceleration Measurement System (MAMS): MAMS
continues to measure the quasi-steady (below 1 Hz) acceleration environment
using the OSS sensor. The High Rate Acceleration Package (HiRAP) telemetry
downlink is enabled to remotely support PFMI operations.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):
Behaving nominally.

Investigating the Structure of Paramagnetic Aggregates from Colloidal
Emulsions (InSPACE): Planned.

Renal Stone: Next in-flight urine collection will occur at the beginning of
next year. The crew was thanked for their continued support for this
experiment with daily pill ingestion at dinnertime.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites
(SHERES): Completed.

Foot (Foot/Ground Reaction Forces during Space Flight): Looking forward to
the new sessions next year.

Materials ISS Experiment (MISSE): In progress. Deployed outside. Nominal and
collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation
(CBOSS-FDI): Looking forward to the next activity.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Students are
looking forward to the next session later this Increment.

Earth Science Toward Exploration Research (ESTER): Looking forward to the
sessions next year.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA): Completed.

Educational Payload Operations (EPO): The ground is looking forward to the
next demo. Crew Earth Observation (CEO): A mosaic of images from the crew's
careful mapping of Lake Titicaca with the 80mm lens last month was published
on Earth Observatory this weekend. In the ongoing review of ISS/CEO imagery,
the ground noted a handsome set of views of the Galapagos and some striking
views with glint over the lower US Great Plains. As for camera data, so far,
investigators have noted a significant reduction in the large time deltas
beginning on 11/28, but suspect that the time is still a bit slow and
getting worse again as of 12/1. Also, many 200mm images are still found a
bit "soft". However, the ground wants to commend and encourage the crew's
continued use of both short lens context views and long lens details of the
same target whenever feasible. This will be particularly helpful when
ISS/CEO begins using the 400mm lens and the doubler for targets requiring
high detail.

Today's CEO targets, in the current LVLH attitude no longer limited by
flight rule constraints on the use of the Lab nadir/science window, except
for some viewing angles, wereSubglacial rivers, SW Libya(400mm-lens: Mapping
swath near or under track for ~30 secs was requested to document sub-glacial
river beds of 450 million years ago [continental ice sheet formed as Africa
slid over the South Pole]. Investigators are trying to reconstruct the
broader, regional patterns of these sinuous, ancient riverbeds. The
scientific interest lies in their preservation and in the highly porous
sediments within them [known to contain water and hydrocarbons]. The ancient
rocks that contain the river beds appear dark in the desert),Subglacial
rivers, W Libya(400mm-lens: Similar mapping swath near or under track for
~30 secs was requested. These ancient riverbeds are known on the ground, but
are not resolvable in available satellite imagery. The ancient rocks that
contain the river beds appear dark in the desert),Chicago, Illinois(nadir
pass. Looking left and right for this urban region on the shore of Lake
Michigan. Also, trying to document lake-effect snow along lake shores,
especially opposite Chicago in downwind Michigan),Baker Island,
Pacific(detail of this remote island and coral reef was requested),La Paz,
Bolivia(nadir pass. Large cloud mass was expected to have shifted),
andPuerto Rico(looking for corals along the coast, and the nearby coastal
urban developments which are causing damage to the corals. Damage is visible
as discoloration [whitening] of the coral).

CEO images can be viewed at the websites.


http://eol.jsc.nasa.gov
http://earthobservatory.nasa.gov

See also the website "Space Station Challenge" at


http://voyager.cet.edu/iss/

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

Elektron O2 generator is powered On, 16A. Vozdukh CO2 scrubber is On (in
Manual Mode 5/3). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed
mode). TCCS (trace contaminant control subsystem) is operating. MCA (major
constituents analyzer) is off (in Life Extending Mode). BMP Harmful
Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS
air conditioner SKV-1 is On; SKV-2 is Off.


SM Working Compartment: Pressure (mmHg) -- 742; temperature (deg C) -- 25.2;
ppO2 (mmHg) -- 154.0; ppCO2 (mmHg) -- 2.8.
SM Transfer Compartment: Pressure (mmHg) -- 765; temperature (deg C) --
20.0.
FGB Cabin: Pressure (mmHg) -- 772; temperature (deg C) -- 22.0.
Node: Pressure (mmHg) -- 766.15; temperature (deg C) -- 23.9 (shell); ppO2
(mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab: Pressure (mmHg) -- 769.19; temperature (deg C) -- 25.0; ppO2
(mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
Joint Airlock (Equip. Lock): Pressure (mmHg) -- 750.07; temperature (deg
C) -- 26.8; shell heater temp (deg C) -- 24.3, ppO2 (mmHg) -- n/a; ppCO2
(mmHg) -- n/a.
PMA-1: Shell heater temp (deg C) -- 24.6
PMA-2: Shell heater temp (deg C) -- 11.7

(n/a = data not available)

Electrical Power Systems (EPS):

Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B
both in Directed Position (non-suntracking, "night glider"/"sun slicer" drag
reduction mode).
SM batteries: Battery #8, formerly known as #7, is still disconnected in
slot #8 for troubleshooting, off-line; all other batteries (7) are in
"Partial Charge" mode.
FGB batteries: Battery #5 is off (capacity restoration mode, ROM); all other
batteries (5) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
INT-2 is operating; INT-1 is Off.
EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is Off; PL-2 MDM is Operational.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1
dropped out 11/22).
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3
dropped out 10/22).

Propulsion System:

Total propellant load available:3680 kg (8113 lb) as of 12/11/03 [SM(755) +
FGB(2573) + Progress M(352) + Progress M-1(0)]. (Capability: SM -- 860 kg;
FGB -- 6120 kg).

Attitude Control Systems:

3 CMGs on-line (CMG-1 failed).
State vector source -- U.S. SIGI-1 (GPS)
Attitude source -- U.S. SIGI-1 (GPS)
Angular rate source -- RGA-1

Flight Attitude:

LVLH YVV (local vertical/local horizontal = "earth-fixed": z-axis in local
vertical, y-axis in velocity vector [yaw: -89.9 deg, pitch: -8.9 deg, roll:
1.8 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally (on string 2).
Ku-band is operating nominally.
Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on
both strings.
MBS: KA power on both strings.
MT: latched and mated at WS4.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is
Off.

ISS Orbit (as of this morning, 7:21am EST [= epoch]):

Mean altitude -- 371.5 km
Apogee -- 375.5 km
Perigee -- 367.5 km
Period -- 92.0 min.
Inclination (to Equator) -- 51.63 deg
Eccentricity -- 0.0005955
Orbits per 24-hr. day -- 15.65
Mean altitude loss in last 24 hours -- 56 m
Revolutions since FGB/Zarya launch (Nov. '98) -- 28935
For more on ISS orbit and worldwide naked-eye visibility dates/times, see
http://www.hq.nasa.gov/osf/station/viewing/issvis.html




--
-------------------

Jacques :-)

Editor: www.spacepatches.info


 




Thread Tools
Display Modes

Posting Rules
You may not post new threads
You may not post replies
You may not post attachments
You may not edit your posts

vB code is On
Smilies are On
[IMG] code is On
HTML code is Off
Forum Jump

Similar Threads
Thread Thread Starter Forum Replies Last Post
ISS Status Report No. 61 - 2003 Jacques van Oene Space Station 0 November 29th 03 05:45 PM
ISS Status Report No. 60 - 2003 Jacques van Oene Space Station 0 November 23rd 03 10:33 AM
ISS Status Report No. 51 - 2003 Jacques van Oene Space Station 0 October 18th 03 10:48 AM
International Space Station Status Report #39 - 2003 Jacques van Oene Space Station 0 August 23rd 03 11:48 AM
Ed Lu Letter from Space #6 Jacques van Oene Space Station 0 July 4th 03 11:10 AM


All times are GMT +1. The time now is 12:12 PM.


Powered by vBulletin® Version 3.6.4
Copyright ©2000 - 2025, Jelsoft Enterprises Ltd.
Copyright ©2004-2025 SpaceBanter.com.
The comments are property of their posters.